OpenStudy (ashy98):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. A.) y - 1 = 4(x - 3) <---? B.) y + 1 = 4(x + 3) C.) y - 3 = 4(x + 1) D.) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y - y1 = m(x-x1) (x1,y1) = given points m = slope

OpenStudy (ashy98):

so i was wrong?

OpenStudy (anonymous):

yeah because points are (-3,-1), so y-(-1) = 4(x-(-3))

OpenStudy (ashy98):

so it would be c?

OpenStudy (anonymous):

y-(-1) =4(x-(-3)) y+1 = 4(x+3) that would be b

OpenStudy (ashy98):

okay can you tell me how you did that ?

OpenStudy (anonymous):

-(-1) = +1

OpenStudy (anonymous):

multiplication of two negative sign is a positive